OpenStudy (anonymous):

If f(x)=-5+1 and g(x)=x^3, what is (g。f)(0)?

OpenStudy (whpalmer4):

Are you sure you have your f(x) correct? I don't see an \(x\) in it.

OpenStudy (anonymous):

Oh sorry yea it's f(x)=-5x+1

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

If memory serves me correctly, that notation is equivalent to g(f(x)), right?

OpenStudy (whpalmer4):

Yes, it is. So, we need to do 1 of 2 things: Either find an algebraic expression for \((g\circ f)(x)\) and evaluate it at \(x = 0\), or evaluate \(u = f(0)\) and then evaluate \(g(u)\)

OpenStudy (anonymous):

Yes I believe so

OpenStudy (whpalmer4):

Let's do both! 2) if \(f(x) = -5x+1\), what does \(f(0) =\)? \(f(0) = -5(0)+1 = \)

OpenStudy (anonymous):

1

OpenStudy (whpalmer4):

Good. Now what is \(g(1) = (1)^3 = \)?

OpenStudy (anonymous):

1

OpenStudy (whpalmer4):

Good. If we had a lot of these to do, it might be worth doing the algebra to find a direct expression for \((g\circ f(x))\). Here's how we would do that. \[f(x) = -5x+1\]\[g(x) = x^3\]Now we take the right hand side of our definition of \(f(x)\) and rewrite our definition of \(g(x)\), replacing \(x\) wherever we encounter it on the right hand side of \(g(x)\) with \(f(x)\). \[g(x) = x^3 = (-5x+1)^3\] We could expand that to \[g(x) = -125x^3+75x^2-15x+1\]but it is probably just as convenient if not more so in the earlier form.

OpenStudy (whpalmer4):

Let's check our work: \[g(0) = (-5(0)+1)^3 = (0+1)^3 = 1^3 = 1\checkmark\] \[g(0) = -125(0)^3 + 75(0)^2 - 15(0)+1 = 1\checkmark\]

OpenStudy (whpalmer4):

Both of the new equations work at \(x=0\), at least. That's a necessary, if not sufficient, condition. We could check another point both ways: \[x = 1,~f(1) = -5(1)+1 = -4, ~g(-4) = (-4)^3 = -64\] \[(g(f(1)) = (-5(1)+1)^3 = (-4)^3 = -64\checkmark\]\[g(f(1)) = -125(1)^3+75(1)^2-15(1)+1 = -125+75-15+1 = -64\checkmark\]So our equations work at 2 different points, I think it is safe to conclude our work is correct.

OpenStudy (whpalmer4):

Any questions about how I combined equations? Or anything else?

OpenStudy (anonymous):

So is our answer 0?

OpenStudy (whpalmer4):

No, our answer was 1. \[g(f(0)) = (-5(0)+1)^3 = (0+1)^3 = 1^3 = 1\] Or \[f(0) = (-5(0)+1) = 0+1 = 1, ~g(f(0)) = g(1) = (1)^3 = 1\]